OpenStudy (anonymous):

Which of the following points does not lie on the line whose equation is x+ y = 7?

OpenStudy (anonymous):

which following

OpenStudy (anonymous):

(-10, 17) (13/3, 8/3) (-1, -6)

OpenStudy (anonymous):

its b

OpenStudy (anonymous):

two negatives make a posotive

OpenStudy (anonymous):

its only b

OpenStudy (anonymous):

ok yes its b thanks....

OpenStudy (anonymous):

no its only b

OpenStudy (anonymous):

k thanks goodbye

OpenStudy (anonymous):

Any point that lies on the line satisfies the equation so those that does not will not lie on it the last point does not satisfu=y the equation and hence does not lie on it

OpenStudy (anonymous):

B satisfies it!! (13/3) +(8/3) = (21/3) = 7

OpenStudy (anonymous):

Negatives donot cancel off each other in addition !!

OpenStudy (anonymous):

two negatives equal a posotive

OpenStudy (anonymous):

isnt it X + Y??

OpenStudy (anonymous):

oh yes

OpenStudy (anonymous):

ima go kill myself bye bye

OpenStudy (anonymous):

-1 + (-6) = -7 not 7

OpenStudy (anonymous):

LOL :P

OpenStudy (anonymous):

bang im dead

OpenStudy (anonymous):

XD chill dude.....

OpenStudy (anonymous):

I was about to mention that, but nacho said the asnwer is b so I was like WTF

OpenStudy (anonymous):

;gjasd;gj